2.1x-6+1.6x=12.5
please help me i am totally lost with this problem!!!!!!

Answer by checkley77(12844)

2.1x-6+1.6x=12.5
3.7x=12.5+6
3.7x=18.5
x=18.5/3.7
x=5 answer.
Proof:
2.1*5-6+1.6*5=12.5
10.5-6+8=12.5
12.5=12.5